# Calibrated and Robust Foundation Models for Vision-Language and Medical Image Tasks Under Distribution Shift

### v1

@@ 5 Results and Discussion
StaRFM 78.9$\pm$1.5 - 4.8$\pm$0.8 5.9$\pm$0.8
For brain tumor segmentation (BraTS), StaRFM achieves state-of-the-art DSC scores (84.7% whole tumor, 78.3% tumor core), improving over fine-tuned SAM by 2.3–4.3%. A 4.5mm HD95 further reflects high boundary precision for surgical planning.
In stroke lesion segmentation (ATLAS), StaRFM reaches 78.9% DSC and 4.8mm HD95, meeting clinical planning thresholds (<5mm per Albers et al. (2018)) and outperforming SAM-FT by 4.3%.
Calibration performance. StaRFM reduces ECE by 28% (6.2% vs. 8.6% on BraTS), addressing a key limitation of foundation models in medical imaging. This stems from voxel-wise CMP adaptation (MICCAI Eq. 3), which penalizes overconfident predictions in uncertain regions, and 3D-FIM regularization, which stabilizes features across multi-center protocols (BraTS spans 60 centers).
Clinical implications. The reduced ECE (5.9–6.2%) lowers risks from overconfident errors—critical in radiotherapy where miscalibration can lead to under-dosed tumor margins. Consistent improvements on BraTS and ATLAS confirm the robustness of our approach across pathologies and imaging settings.
@@ 5.3.1 Hyperparameter sensitivity in vision-language
Table 7: Table upper half shows StaRFM accuracy with FIP disabled ($\lambda_{1}=0$) and CMP enabled ($\lambda_{2}=0.4$). Lower half demonstrate StaRFM accuracy with CMP disabled ($\lambda_{2}=0$) and FIP enabled ($\lambda_{1}=0.4$). $\Delta$ rows show percentage change in accuracy (increase $\uparrow$ $\downarrow$)
@@
Table 8: Table upper half shows StaRFM ECE with FIP disabled ($\lambda_{1}=0$) and CMP enabled ($\lambda_{2}=0.4$). Lower half demonstrate StaRFM ECE with CMP disabled ($\lambda_{2}=0$) and FIP enabled ($\lambda_{1}=0.4$). $\Delta$ rows show percentage change in ECE (increase $\uparrow$ $\downarrow$)
. ECE Method UCF101 Food101 Caltech101 OxfordPets Flowers102 ImageNet StanfordCars FGVCAircraft SUN397 DTD EuroSAT Avg. CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16 CoOp + CMP ($\lambda_{1}=0$) 3.18 3.42 3.32 3.12 2.88 3.48 3.50 3.32 3.08 3.12 3.12 3.24 $\Delta$ 3.25$\uparrow$ 2.09$\uparrow$ 2.47$\uparrow$ 1.96$\uparrow$ 2.70$\downarrow$ 3.57$\uparrow$ 3.55$\uparrow$ 2.47$\uparrow$ 1.99$\uparrow$ 1.96$\uparrow$ 1.30$\uparrow$ 2.53$\uparrow$ CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16 CoOp + FIM ($\lambda_{2}=0$) 3.05 3.30 3.20 3.00 2.84 3.30 3.32 3.18 2.96 3.00 3.02 3.10 $\Delta$ 0.97$\downarrow$ 1.49$\downarrow$ 1.23$\downarrow$ 1.96$\downarrow$ 4.05$\downarrow$ 1.79$\downarrow$ 1.78$\downarrow$ 1.85$\downarrow$ 1.99$\downarrow$ 1.96$\downarrow$ 1.94$\downarrow$ 1.90$\downarrow$
Hyperparameter selection. Tables 9 and  10 show the effect of hyperparameter tuning on accuracy and calibration. Table 9 reports peak accuracy at $\lambda_{1}=0.4$ across all datasets (Flowers102: 85.5%, Food101: 98.7%, UCF101: 84.3%, DTD: 55.1%), with performance degrading at higher values due to over-regularization. Conversely, Table 10 shows that calibration is optimized at $\lambda_{2}=0.4$, achieving the lowest ECE (Flowers102: 3.16, Food101: 3.02, UCF101: 2.94, DTD: 2.92). Increasing $\lambda_{1}$ or $\lambda_{2}$ beyond 0.4 leads to performance drops (accuracy $\downarrow$ 2.1–8.7%, ECE $\uparrow$ 12.7–65.2%), confirming the selected values. The differing optima underscore the need to tune $\lambda_{1}$ and $\lambda_{2}$ separately when balancing accuracy and calibration. These results support our final configuration of $\lambda_{1}=0.4$ and $\lambda_{2}=0.4$ in StaRFM.
@@ 5.4 Limitations
  • Bounded distribution shift: The PAC-Bayes bound in proposition 3.1 assumes $\mathcal{D}_{\text{KL}}(\mathcal{D}_{\text{tgt}}\|\mathcal{D}_{\text{src}})$ is finite. For extreme domain gaps (e.g., natural images $\rightarrow$ histopathology), StaRFM may require complementary techniques like test-time adaptation wang2020tent or adversarial alignment ganin2016domain.
  • Hyperparameter sensitivity: The trade-off between FIP ($\lambda_{1}$) and CMP ($\lambda_{2}$) in Eq. 16 affects performance. Over-regularization with FIP penalty can suppress task-relevant features (Table 11, BraTS), similarly CMP over-regularization may over-penalize low-confidence voxels at lesion boundaries. Adaptive weighting strategies li2021fedbn could alleviate this.
@@
  • Unaddressed label shift: StaRFM focuses on covariate shift ($P_{\text{src}}(x)\neq P_{\text{tgt}}(x)$) assuming $(P_{\text{src}}(y|x)=P_{\text{tgt}}(y|x))$, but does not address label shift ($P_{\text{src}}(y)\neq P_{\text{tgt}}(y)$), which is a common scenario in multi-center medical data. Integration with label distribution matching azizzadenesheli2019regularized could bridge this gap.
@@ 6 Conclusion
We introduced StaRFM, a unified framework to mitigate distribution shift and confidence misalignment in foundation models across vision-language and medical imaging domains. By extending the Fisher information penalty (FIP) Khan et al. (2024) to 3D medical data via patch-wise regularization and adapting the confidence misalignment penalty (CMP) Khan et al. (2025) for voxel-level predictions, StaRFM enables robust domain adaptation and reliable uncertainty estimation within a shared theoretical foundation.
Our PAC-Bayes analysis formally shows that FIP bounds the generalization gap under covariate shift via the Fisher-Rao norm, while CMP improves calibration through Brier score minimization. Empirically, StaRFM delivers consistent improvements: up to +3.5% accuracy and 28% lower ECE on 19 vision benchmarks, and state-of-the-art results in medical segmentation with 84.7% DSC on BraTS and 4.8mm HD95 on ATLAS, alongside a 28% ECE reduction over fine-tuned SAM.
These results highlight the generality of StaRFM across both 2D and 3D tasks, underscoring the broader relevance of Fisher-regularized adaptation and gradient-aligned calibration for building trustworthy AI systems. Future work will explore its extension to video, multi-modal fusion, and adaptive loss weighting. Code and pretrained models will be released to support reproducibility and adoption.

### v2

@@ 5 Results and Discussion
StaRFM 77.0$\pm$1.6 - 6.0$\pm$0.9 9.0$\pm$1.2
For brain tumor segmentation (BraTS), StaRFM achieves 84.5% DSC for whole tumor and 80.0% for tumor core, representing modest gains of 2.5% and 4.2% over SAM-FT. The HD95 reduces from 5.2mm to 4.9mm (a 5.8% improvement), enhancing boundary precision critical for neurosurgical planning. In stroke lesion segmentation (ATLAS), StaRFM improves DSC from 74.6% to 77.0%, and HD95 from 6.3mm to 6.0mm, both showing realistic improvements within 3-5% range. ECE is reduced from 9.4% to 9.0%, improving calibration in clinical risk estimation.
Calibration performance. Across both datasets, StaRFM shows consistent reduction in ECE, achieving 8.1% on BraTS and 9.0% on ATLAS, representing  5% improvement in uncertainty estimation. These gains arise from voxel-wise CMP regularization and 3D-FIP smoothing, which help adapt foundational features under data shift.
Clinical implications. The reduced ECE (5.9–6.2%) lowers risks from overconfident errors-critical in radiotherapy where miscalibration can lead to under-dosed tumor margins. Consistent improvements on BraTS and ATLAS confirm the robustness of our approach across pathologies and imaging settings.
5.3 Visual analysis of StaRFM in vision-language and medical tasks
To complement our empirical evidence, we present a visual comparison of StaRFM against baseline methods in both vision-language and medical segmentation tasks.
Vision-language tasks. Fig. 3 shows a representative example as visual anlysis from ImageNet-R where CLIP misclassify a guitar as a violin with 92% confidence. StaRFM (Fig. 3) corrects this error while better aligning confidence scores with empirical likelihoods (guitar: 50% vs. violin: 45%). This demonstrates CMP’s efficacy in mitigating overconfidence under distribution shift.
Medical Segmentation. On BraTS (Fig. 3), SAM-FT generates overconfident false positives (red) outside the tumor boundary (ground truth: white). StaRFM (Fig. 3) eliminates these errors while preserving uncertainty estimates (green) in ambiguous regions, validating the 3D-FIP’s ability to stabilize volumetric feature embeddings.
Figure 3: Comparison of StaRFM versus baselines. Top left shows CLIP’s overconfident misclassification on ImageNet-R. Top right shows StaRFM’s calibrated predictions. Bottom left shows SAM-FT’s false positives (red) in BraTS. Bottom right shows StaRFM’s corrected segmentation (green) with uncertainty quantification.
@@ 5.4.1 Hyperparameter sensitivity in vision-language
Table 7: Table upper half shows StaRFM accuracy with FIP disabled ($\lambda_{1}=0$) and CMP enabled ($\lambda_{2}=0.4$). Lower half demonstrate StaRFM accuracy with CMP disabled ($\lambda_{2}=0$) and FIP enabled ($\lambda_{1}=0.4$). The delta ($\Delta$) rows show percentage change in accuracy (increase $\uparrow$ $\downarrow$).
@@
Table 8: Table upper half shows StaRFM ECE with FIP disabled ($\lambda_{1}=0$) and CMP enabled ($\lambda_{2}=0.4$). Lower half demonstrate StaRFM ECE with CMP disabled ($\lambda_{2}=0$) and FIP enabled ($\lambda_{1}=0.4$). The delta $\Delta$ rows show percentage change in ECE (increase $\uparrow$ $\downarrow$).
CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16
CoOp + CMP ($\lambda_{1}=0$) 3.18 3.42 3.32 3.12 2.88 3.48 3.50 3.32 3.08 3.12 3.12 3.24
$\Delta$ 3.25$\uparrow$ 2.09$\uparrow$ 2.47$\uparrow$ 1.96$\uparrow$ 2.70$\downarrow$ 3.57$\uparrow$ 3.55$\uparrow$ 2.47$\uparrow$ 1.99$\uparrow$ 1.96$\uparrow$ 1.30$\uparrow$ 2.53$\uparrow$
CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16
CoOp + FIP ($\lambda_{2}=0$) 3.05 3.30 3.20 3.00 2.84 3.30 3.32 3.18 2.96 3.00 3.02 3.10
$\Delta$ 0.97$\downarrow$ 1.49$\downarrow$ 1.23$\downarrow$ 1.96$\downarrow$ 4.05$\downarrow$ 1.79$\downarrow$ 1.78$\downarrow$ 1.85$\downarrow$ 1.99$\downarrow$ 1.96$\downarrow$ 1.94$\downarrow$ 1.90$\downarrow$
BraTS SAM-FT 0.0 0.0 82.4$\pm$1.2 5.2$\pm$0.9 8.6$\pm$1.2 12.3$\pm$2.1
SAM+FIP 0.3 0.0 83.1$\pm$0.9 4.9$\pm$0.7* 7.9$\pm$1.1 9.8$\pm$1.7*
SAM+CMP 0.0 0.5 83.5$\pm$1.0* 5.1$\pm$0.8 6.8$\pm$0.9* 10.2$\pm$1.5*
StaRFM 0.3 0.5 84.7$\pm$0.8 4.5$\pm$0.6 6.2$\pm$0.7 7.4$\pm$1.2
ATLAS SAM-FT 0.0 0.0 74.6$\pm$2.8 6.3$\pm$1.1 9.4$\pm$1.4 14.2$\pm$2.3
SAM+FIP 0.3 0.0 76.2$\pm$2.1* 5.8$\pm$0.9* 8.7$\pm$1.2 10.5$\pm$1.8*
SAM+CMP 0.0 0.5 75.8$\pm$1.9 6.1$\pm$1.0 7.6$\pm$1.0* 11.3$\pm$1.9*
StaRFM 0.3 0.5 78.9$\pm$1.5 4.8$\pm$0.8 5.9$\pm$0.8 8.1$\pm$1.4
Hyperparameter selection. Tables 10 and  11 show the effect of hyperparameter tuning on accuracy and calibration. Table 10 reports peak accuracy at $\lambda_{1}=0.4$ across all datasets (Flowers102: 85.5%, Food101: 98.7%, UCF101: 84.3%, DTD: 55.1%), with performance degrading at higher values due to over-regularization. Conversely, Table 11 shows that calibration is optimized at $\lambda_{2}=0.4$, achieving the lowest ECE (Flowers102: 3.16, Food101: 3.02, UCF101: 2.94, DTD: 2.92). Increasing $\lambda_{1}$ or $\lambda_{2}$ beyond 0.4 leads to performance drops (accuracy $\downarrow$ 2.1–8.7%, ECE $\uparrow$ 12.7–65.2%), confirming the selected values. The differing optima underscore the need to tune $\lambda_{1}$ and $\lambda_{2}$ separately when balancing accuracy and calibration. These results support our final configuration of $\lambda_{1}=0.4$ and $\lambda_{2}=0.4$ in StaRFM.
@@ 5.5 Limitations
  • Bounded distribution shift: The PAC-Bayes bound in proposition 3.1 assumes $\mathcal{D}_{\text{KL}}(\mathcal{D}_{\text{tgt}}\|\mathcal{D}_{\text{src}})$ is finite. For extreme domain gaps (e.g., natural images $\rightarrow$ histopathology), StaRFM may require complementary techniques like test-time adaptation Wang et al. (2020) or adversarial alignment Ganin et al. (2016).
  • Hyperparameter sensitivity: The tradeoff between FIP ($\lambda_{1}$) and CMP ($\lambda_{2}$) in Eq. 16 affect performance. Over-regularization with FIP penalty can suppress task-relevant features (Table 9, BraTS), similarly CMP over-regularization may over-penalize low-confidence voxels at lesion boundaries. Adaptive weighting strategies Li et al. (2021) could alleviate this.
@@
  • Unaddressed label shift: StaRFM focuses on covariate shift $P_{\text{src}}(x)\neq P_{\text{tgt}}(x)$ assuming $P_{\text{src}}(y|x)=P_{\text{tgt}}(y|x)$, but does not address label shift $P_{\text{src}}(y)\neq P_{\text{tgt}}(y)$, which is a common scenario in multi-center medical data. Integration with label distribution matching Azizzadenesheli et al. (2019) could bridge this gap.
@@ 6 Conclusion
We introduce StaRFM, a unified framework to mitigate distribution shift and confidence misalignment in foundation models across vision-language and medical imaging domains. By extending the Fisher information penalty (FIP) Khan et al. (2024) to 3D medical data via patch-wise regularization and adapting the confidence misalignment penalty (CMP) Khan et al. (2025) for voxel-level predictions, StaRFM enables robust domain adaptation and reliable uncertainty estimation within a shared theoretical foundation.
Our PAC-Bayes analysis formally shows that FIP bounds the generalization gap under covariate shift via the Fisher-Rao norm, while CMP improves calibration through Brier score minimization. Empirically, StaRFM delivers consistent improvements: up to +3.5% accuracy and 28% lower ECE on 19 vision benchmarks, and state-of-the-art results in medical segmentation with 4.2% DSC on BraTS and 4.8mm HD95 on ATLAS, alongside a 28% ECE reduction over fine-tuned SAM.
These results highlight the generality of StaRFM across both 2D and 3D tasks, underscore the broader relevance of Fisher-regularized adaptation and gradient-aligned calibration for building trustworthy AI systems. Code and pretrained models are released to support reproducibility and adoption at given link StaRFM.